Circular No. 7365

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-3 M. L. McCollough and G. J. Fishman, Marshall Space Flight Cen-

ter, for the Compton Gamma Ray Observatory (CGRO) BATSE team; and E. B. Waltman, Naval Research Laboratory, report: “The galactic jet source Cygnus X-3 has been exhibiting very low flux levels in radio and hard x-rays (20–100 keV) since Feb. 4. Such behavior usually precedes major radio flaring episodes, with radio fluxes > 3 Jy (McCollough et al. 1999, Ap.J. 517, 951). Very low radio fluxes of < 30 mJy have been observed at 2 GHz with the Green Bank Interferometer between Feb. 4 and 9. Daily averages for Jan. 29–Feb. 11 of the hard-x-ray flux, measured using the BATSE/CGRO earth-occultation technique, have been below the 1-day detection limit of ∼ 100 mCrab. Previous observations of Cyg X-3 with periods of very-low-radio and hard-x-ray flux, having durations from days to a few weeks, are followed by major radio flares. The last such episode occurred in 1997 (IAUC 6673). Target-of-opportunity observations have been initiated for both RXTE and CGRO. Observations at other wavelengths are encouraged during the next few weeks.”

SW URSAE MAJORIS J. A. Mattei, AAVSO, reports that this cataclysmic variable is in out-

burst for the first time since its 20-day superoutburst in Oct. 1997 (IAUC 6760), when the AAVSO International Database shows that the star reached mean mag 10.7 at maximum. R. Zissell, South Hadley, MA, found SW UMa at CCD V = 16.55 on Feb. 9.192 UT. Selected visual magnitude estimates forwarded by Mattei: Feb. 11.243, [15.6 (G. Hanson, Cave Creek, AZ); 11.926, 14.1 (C. Jones, Laindon, Essex, England); 12.028, 12.4 (P. Schmeer, Saarbruecken, Germany); 12.126, 11.5 (S. Tracy, North Granby, CT); 12.74, 10.7 (J. Gunther, Peyrins, France); 13.060, 10.5 (J. Bortle, Stormville, NY); 14.044, 10.5 (G. Poyner, West Midlands, England).

SUPERNOVA 1999gq IN NGC 4523 Unfiltered CCD magnitudes: 1999 Dec. 25.806 UT, 15.1 (S. Yoshida

and K. Kadota, Ageo, Saitama, Japan); 2000 Jan. 21.191, 14.97 (M. Suárez, Las Palmas, Canary Islands); 28.701, 15.0 (Yoshida and Kadota).

2000 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green